Lucas has 37 coins, all nickels and dimes in his piggy bank. The value of the coins is $3.10. How many dimes does Lucas have?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52787)

.
Lucas has 37 coins, all nickels and dimes in his piggy bank.
The value of the coins is $3.10. How many dimes does Lucas have?
~~~~~~~~~~~~~~~~~~~

D dimes;  (37-D) nickels.


Write the total money equation

    10*D + 5*(37-D) = 310  cents.


Solve it and find D

    10*D + 5*37 - 5*D = 310

    10*D - 5*D = 310 - 5*37

        5*D   =     125

          D   =     125/5 = 25.


ANSWER.  25 dimes.

Solved.



Answer by greenestamps(13200)



The typical formal algebraic solution shown by the other tutor is of course fine; you should understand it and be able to solve similar problems by that method.

For a mental solution method....

Note that if all 37 coins were dimes the total value would be $3.70.
That is $0.60 more than the actual total of $3.10.
Since each nickel is worth $0.10-$0.05 = $0.05 less than each dime, the number of nickels must be $0.60/$0.05 = 60/5 = 12.
Therefore the number of dimes is 37-12 = 25

ANSWER: 25

Note this method can be used to solve a wide range of similar problems quickly, without formal algebra.